OpenStudy (anonymous):

What is the solution to –4|–2x + 6| = –24?

OpenStudy (solomonzelman):

\(1.\) Divide by -4 on both sides.

OpenStudy (solomonzelman):

what do you get after dividing by -4 ?

OpenStudy (solomonzelman):

i didn't mean to interrupt you, always go ahead and say whatever you have to say. Alays welcome... please:)

OpenStudy (solomonzelman):

all i was going to post is this abstract example --------------------------------------- a|cx+b|=f |cx+b|=(f/a) cx+b=(f/a) or cx+b=-(f/a) cx=(f/a)-b cx=-(f/a)-b x=((f/a)/c)-b/c x=-((f/a)/c)-b/c ---------------------------------------

OpenStudy (solomonzelman):

but, we can do it together if you wish to proceed. Please divie by -4 on both sides and tell me what you get.